# Zero-shot Concept Bottleneck Models

### v1

@@ 5.1 Settings
Datasets. We used 12 classification datasets of various image domains: Aircraft (Air) (Maji et al., 2013), Bird (Welinder et al., 2010), Caltech-101 (Cal) (Fei-Fei et al., 2004) Car (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Euro) (Helber et al., 2019), Flower (Flo) (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), ImageNet (IN) (Russakovsky et al., 2015), Pet (Parkhi et al., 2012), SUN397 (Xiao et al., 2010), and UCF-101 (Soomro, 2012). They are often used to evaluate the zero-shot generalization performance of VLMs (Radford et al., 2021; Zhou et al., 2022). In the training head scenario, we randomly split a training dataset into $9:1$ and used the former as the training set and the latter as the validation set. For ImageNet, we set the split ratio $99:1$.
Zero-shot Baselines. Since no zero-shot baselines of CBMs exist, we compare our Z-CBMs with the zero-shot inference of a black-box VLM and variants of Z-CBMs in terms of regression algorithms and concept banks. For more details, please see Appendix B.
@@
Training Head Baselines. To compare Z-CBMs with existing vision-language-based CBMs, we evaluated models in a relaxed setting where the models are trained on target datasets. In this setting, we applied Z-CBMs to linear probing of VLMs, i.e., fine-tuning only a linear head layer on the feature extractors of VLMs; we refer to this pattern LP-Z-CBM. As the baselines, we used Lable-free CBM (Oikarinen et al., 2023), LaBo (Yang et al., 2023), and CDM (Panousis et al., 2023). We performed these methods based on their publicly available code repositories.
@@
Evaluation Metrics. For evaluating predicted concepts, we measured CLIP-Score (Radford et al., 2021; Hessel et al., 2021), which is the cosine similarity between image and text embeddings on CLIP, i.e., higher is better. CLIP-Score between input images and concepts intuitively indicates how well the predicted concept explains the image. Thus, it performs as an indicator to evaluate the quality of the input-to-concept inference. Concretely, we measured averaged CLIP-Scores between test images and the predicted concept texts, where we extracted the top 10 concepts from sorted concepts in descending order by absolute concept importance scores for each model. Note that, to compute CLIP-Score, we used CLIP ViT-L/14, which is a different pre-trained model from CLIP ViT-B/32. Furthermore, we used concept coverage to evaluate the Z-CBM’s predicted concepts. Concept coverage |{${|\{c^{\mathrm{Z}}_{i}\}\cap\{c^{\mathrm{R}}_{i}\}|}/{|\{c^{\mathrm{R}}_{i}\}|}$ is the ratio of overlap between Z-CBM’s concepts with non-zero coefficients $\{c^{\mathrm{Z}}_{i}\}\subset C$ and reference concepts $\{c^{\mathrm{R}}_{i}\}\subset C$ predicted by vision-language-based CBMs that require training. This metric evaluates the extent to which the Z-CBM yields concepts that are close to those derived in the target training when using the shared concept bank $C$. Specifically, we computed the average concept coverage across test samples by using the GPT-generated concept banks by (Oikarinen et al., 2023), and reference concepts of Label-free CBMs; we used concepts with contribution scores greater than 0.05 as $\{c^{\mathrm{R}}_{i}\}$ by following (Oikarinen et al., 2023). We also report top-1 test accuracy as the target classification task performance.
@@ 5.2 Quantitative Evaluation of Predicted Concepts
Table 1: CLIP-Score on 12 classification datasets. We compute the averaged CLIP-Scores between images and concepts with top-10 absolute coefficients. Complete results appear in Table 7.
@@
Label-free CBM 0.6982
LaBo 0.7078
CDM 0.7141
Z-CBM (ALL) 0.7754
Table 2: Concept coverage (%) of Z-CBMs on 12 classification datasets. Complete results appear in Table 8.
@@ 5.3 Evaluation of Human Intervention
Human intervention in the output concept is an essential feature shared by the CBM family for debugging models and modifying the output concepts to make the final prediction accurate. Here, we evaluate the reliability of Z-CBMs through two types of intervention: (i) concept deletion and (ii) concept insertion. In concept deletion, we confirm the dependence on the predicted concepts by removing the concept with non-zero coefficients in ascending, descending, and random orders. Fig. 3 is the results on Bird by varying the deletion ratio. The accuracy of Z-CBMs significantly dropped with the smaller deletion ratio in the case of descent. This indicates that Z-CBM accurately selects the important concepts through concept regression and predicts the final label based on the concepts. In the case of ascent, the accuracy slowly and steadily decreases, suggesting that the Z-CBMs are not biased toward limited concepts and that all of the selected concepts are essential.
Figure 3: Concept Deletion (Bird)
Figure 4: Concept Insertion (Bird)
In concept insertion, we add ground truth concepts to the predicted concepts with non-zero coefficients and then re-compute concept regression on the intervened concept set. Specifically, we used linear regression as the algorithm in concept regression and then predicted target labels by the weighted averaged intervened concept vectors by Eq. (4). As the ground truth concepts, we used the fine-grained multi-labels annotated for Bird (Welinder et al., 2010). Fig. 4 demonstrates the top-1 accuracy of the intervened Z-CBMs. The performance improved as the number of inserted concepts per sample increased. This indicates that Z-CBMs can correct the final output by modifying the concept of interest through intervention.
Figure 5: Qualitative evaluation of predicted concepts on the ImageNet validation set. While Label-free CBMs sometimes hallucinate invisible concepts or ignore important concepts, Z-CBMs with lasso consistently provide realistic and dominant concepts in input images with diverse vocabulary. NOT prefix denotes that the concept has negative coefficients.
@@ 5.4 Qualitative Evaluation of Predicted Concepts
We demonstrate the qualitative evaluation of predicted concepts by Label-free CBMs and Z-CBMs when inputting the ImageNet validation examples in Fig. 5; we also show the results of Z-CBMs using linear regression to compute the importance coefficients instead of lasso. Overall, Z-CBMs tend to accurately predict realistic and dominant concepts that appear in input images even though they are not trained on target tasks. For instance, in the first row, Z-CBM predicts various concepts related to dogs, clothes, and background, whereas Label-free CBM focuses on clothes and ignores dogs and background. This difference may be caused by the fact that the image-to-concept mapping of Z-CBMs is not biased toward the label information because it does not train on the target data. Conversely, like the second row, Z-CBMs tend to concentrate on global regions and miss the concepts in local regions; this can be alleviated by intervening in the concept prediction (see Sec. 5.3).
For the comparison of linear regression and lasso, we can see that Z-CBM (Linear Reg.) tends to produce concepts that are related to each other. In fact, quantitatively, we also found that the averaged inner CLIP-Scores among the top-10 concepts of lasso (0.6855) is significantly lower than that of linear regression (0.7826). These results emphasize the advantage of using sparse modeling in concept regression to select mutually exclusive concepts based on the concept bank containing abundant vocabulary.
@@ 5.6.2 Effects of Concept Bank
Table 3: Top-1 accuracy on 12 classification datasets with CLIP ViT-B/32. Complete results appear in Table 9.
@@
Backbone VLM Top-1 Acc. Top-1 Acc. CLIP-Score
@@
CLIP ViT-B/32 61.88 62.70 0.7766
CLIP ViT-L/14 72.87 73.19 0.7881
OpenCLIP ViT-H/14 77.20 77.81 0.7910
OpenCLIP ViT-G/14 79.03 78.27 0.8095
SigLIP ViT-SO400M/14 82.27 81.74 0.8241
DFN ViT-H/14 83.85 83.40 0.8337
5.6.3 Effects of Concept Regressor
Z-CBMs allow users to choose arbitrary sparse linear regression algorithms according to their demands, as discussed in Sec. 4. Here, we compare the performance of Z-CBMs with multiple sparse linear regression algorithms: lasso (Tibshirani, 1996), elastic net (Zou & Hastie, 2005), and sparsity-constrained optimization with HTP (Yuan et al., 2014). Further, we evaluate these sparse algorithms by comparing them with non-sparse algorithms to compute the importance of concepts: CLIP Similarity, which uses the cosine similarity computed on CLIP as the importance, and linear regression. Table 6 shows the performance, where sparsity is a ratio of non-zero importance coefficients to the total number of concept candidates. While the sparse linear regression algorithms achieved top-1 accuracy scores at the same level, the non-sparse algorithms failed to accurately predict labels from importance-weighted concepts. Additionally, linear regression has unstable numerical computation due to the rank-deficient of the Gram matrix of $F_{C_{x}}$ when the feature dimension $d$ is smaller than the concept retrieval size $K$. In contrast, lasso can avoid this by sparse regularization. These results indicate that the concept selection by sparse linear regression is crucial in Z-CBMs. In this sense, we can interpret our concept regression as a re-ranking method of the CLIP similarity. Elastic net was the best in accuracy, but it selected more concepts than the other sparse algorithms. This is because elastic net selects all highly correlated concepts to derive a unique solution by combining $\ell_{1}$ and $\ell_{2}$ regularization (Hastie et al., 2015). HTP explicitly limits the number of concepts selected to 256, so while it achieves the highest sparsity, it had the lowest accuracy of the sparse algorithms due to the shortage of concepts for explanation.
@@ 5.6.3 Effects of Concept Regressor
Concept Bank Vocab. Size Top-1 Acc. CLIP-Score
@@
Label-free CBM w/ GPT-3 (ImageNet Class) 4K 58.00 0.7056
CDM w/ GPT-3 (ImageNet Class) 4K 62.52 0.7445
GPT-3 (ImageNet Class) 4K 59.18 0.6276
Noun Phrase (Flickr30K) 45K 61.52 0.6770
Noun Phrase (CC3M) 186K 62.38 0.7109
Noun Phrase (CC12M) 2.58M 62.42 0.7671
Noun Phrase (YFCC15M) 2.20M 62.45 0.7679
Noun Phrase (ALL) 5.12M 62.70 0.7746
Table 6: Performance of Z-CBMs varying concept regressor on ImageNet with CLIP ViT-B/32.
Concept Regressor Top-1 Acc. Sparsity CLIP-Score
CLIP Similarity 14.66 0.0000 0.8117
Linear Regression 52.88 0.0000 0.7076
Lasso 62.70 0.8201 0.7746
Elastic Net 62.84 0.7311 0.7818
Sparsity-Constrained (HTP) 62.54 0.8750 0.7795
@@ 6 Conclusion
Impact Statements.
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ C.1 Detailed Results for All Datasets
Table 7, 8, and 9 shows all of the results on the 12 datasets omitted in Table 1, 2, and 3, respectively.

### v2

@@ 5.1 Settings
Datasets. By following previous studies (Radford et al., 2021), we evaluated Z-CBMs on 12 diverse image classification datasets: Aircraft (Air) (Maji et al., 2013), Bird (Welinder et al., 2010), Caltech-101 (Cal) (Fei-Fei et al., 2004) Car (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Euro) (Helber et al., 2019), Flower (Flo) (Nilsback & Zisserman, 2008), Food (Bossard et al., 2014), ImageNet (IN) (Russakovsky et al., 2015), Pet (Parkhi et al., 2012), SUN397 (Xiao et al., 2010), and UCF-101 (Soomro, 2012). In the training head scenario, we randomly split a training dataset into $9:1$ and used the former as the training set and the latter as the validation set. For ImageNet, we set the split ratio $99:1$.
Zero-shot Baselines. For the zero-shot baseline, our Z-CBMs with the zero-shot inference of a black-box VLM and ConSe (Norouzi et al., 2014), which is a zero-shot classification method predicting a class label with a weighted sum of ImageNet concept features. For more details, please see Appendix B.
@@
Training Head Baselines. To compare Z-CBMs with existing VLM-based CBMs, we evaluated models trained on target datasets. In this setting, Z-CBMs were applied to linear probing of VLMs, i.e., fine-tuning only a linear head layer on the feature extractors of VLMs; we refer to this pattern LP-Z-CBM. As the baselines, we used Lable-free CBM (Oikarinen et al., 2023), LaBo (Yang et al., 2023), and CDM (Panousis et al., 2023). We performed these methods based on their repositories.
@@
Evaluation Metrics. To evaluate predicted concepts, we used the SigLIP-Score, the cosine similarity between image and text embeddings on SigLIP (Zhai et al., 2023) (higher is better). This score indicates how well a predicted concept explains an image (Radford et al., 2021; Hessel et al., 2021), serving as a quality indicator for input-to-concept inference. Specifically, we averaged SigLIP-Scores between test images and their predicted concept texts for the top 10 concepts, ranked by absolute importance scores. Concretely, we measured the average SigLIP-Scores between test images and the predicted concept texts, where we extracted the top 10 concepts from sorted concepts in descending order by absolute concept importance scores for each model. We also used concept recall to evaluate Z-CBM’s predicted concepts. Top-$K$ concept recall ${|C^{\mathrm{Z}}\cap C^{\mathrm{R}}|}/{K}$ measures the overlap between Z-CBM’s top-$K$ concepts $C^{\mathrm{Z}}=\{c^{\mathrm{Z}}_{i}\}^{K}_{i=1}\subset\mathbf{C}$ (with non-zero coefficients) and $N_{\mathrm{R}}$ reference concepts $C^{\mathrm{R}}=\{c^{\mathrm{R}}_{i}\}^{N_{\mathrm{R}}}_{i=1}\subset\mathbf{C}$ from VLM-based CBMs requiring training. This metric assesses concept overlap between Z-CBMs and a training-based CBM using the same concept bank $\mathbf{C}$, indicating Z-CBM’s approximation of a trained model’s concepts. Specifically, we averaged concept recall at $K=10$ across test samples, using the GPT-generated concept banks (Oikarinen et al., 2023), and reference concepts of Label-free CBMs Following (Oikarinen et al., 2023), $C^{\mathrm{R}}$ comprised concepts with contribution scores >0.05>0.05. Finally, we report top-1 test accuracy for target classification performance.
@@ 5.2 Quantitative Evaluation of Predicted Concepts
Table 2: SigLIP-Score
@@
Label-free CBM 0.5485
LaBo 0.5419
CDM 0.5714
Z-CBM (ALL) 0.6309
Table 3: Concept Recall (%)
@@ 5.3 Evaluation of Human Intervention
Human intervention in concepts is an essential feature shared by the CBM family for debugging models and modifying the output to make the final prediction accurate. In addition to interventions in existing concepts in the concept bank, Z-CBMs allows interventions in arbitrary concepts described in natural language. We evaluate Z-CBMs via two types of intervention: (i) concept deletion and (ii) concept insertion. In concept deletion, we confirm the dependence on the predicted concepts by removing the concept with non-zero coefficients in ascending, descending, and random orders. Fig. 4 shows the results on Bird by varying the deletion ratio. The accuracy of Z-CBMs largely dropped with the smaller deletion ratio in the descent cases, indicating that Z-CBM selects the important concepts via concept regression and relies on them for the final prediction. In the ascent cases, the accuracy slowly and steadily decreases, suggesting that the Z-CBMs are not biased toward limited concepts and that all of the selected concepts are essential.
In concept insertion, we first predict concepts by concept regression and add randomly selected ground-truth concepts to the output non-zero concept set. Then, we re-run concept regression with linear regression on this modified concept set and predict the final label prediction by Eq. (5). As the ground truth concepts, we used the attribute labels of Bird (Welinder et al., 2010). Fig. 4 shows the top-1 accuracy of the intervened Z-CBMs. In addition to random selection, we performed a sophisticated intervention method called ECTP (Shin et al., 2023). The performance improved as the number of inserted concepts per sample increased for both cases. This indicates that Z-CBMs can correct the final output by modifying the concept of interest through intervention.
Figure 3: Concept Deletion
Figure 4: Concept Insertion
@@ 5.4 Qualitative Evaluation of Predicted Concepts
We demonstrate the qualitative evaluation of predicted concepts by Label-free CBMs, Z-CBMs with Liner Reg., and Z-CBMs with Lasso when inputting the ImageNet validation examples in Fig. 5. Overall, Z-CBMs tend to accurately predict realistic and dominant concepts that appear in input images, even though they are not trained on target tasks. For instance, Z-CBM predicts various concepts related to dogs, clothes, and background, whereas Label-free CBM focuses on clothes and ignores dogs and background. This difference may be caused by the fact that the image-to-concept mapping of Z-CBMs is not biased toward the label information because it does not train on the target data. For the comparison of linear regression and lasso, Z-CBM (Linear Reg.) tends to produce semantically overlapped concepts. In fact, quantitatively, we also found that the averaged inner SigLIP-Scores among the top-10 concepts of lasso (0.5552) is significantly lower than that of linear regression (0.7425). These results emphasize the advantage of using sparse modeling in concept regression to select mutually exclusive concepts from the large concept bank.
Figure 5: Qualitative evaluation of predicted concepts on the ImageNet validation set. While Label-free CBMs sometimes hallucinate invisible concepts or ignore important concepts, Z-CBMs with lasso provide realistic and dominant concepts in inputs with diverse vocabulary. NOT prefix denotes that the concept has negative coefficients similar to (Oikarinen et al., 2023) and (Panousis et al., 2023).
@@ 5.6.2 Effects of Concept Bank
Table 4: Top-1 accuracy on 12 classification datasets with CLIP ViT-B/32. Complete results appear in Appendix D-A.
@@
Backbone VLM Top-1 Acc. Top-1 Acc. SigLIP-Score
@@
CLIP ViT-B/32 61.88 62.70 0.6498
CLIP ViT-L/14 72.87 73.19 0.6608
OpenCLIP ViT-H/14 77.20 77.81 0.6790
OpenCLIP ViT-G/14 79.03 78.27 0.6810
DFN ViT-H/14 83.85 83.40 0.7038
@@
Concept Bank Vocab. Size Top-1 Acc. SigLIP-Score
@@
Label-free CBM w/ GPT-3 (ImageNet Class) 4K 58.00 0.5896
CDM w/ GPT-3 (ImageNet Class) 4K 62.52 0.6193
GPT-3 (ImageNet Class) 4K 59.18 0.5407
Noun Phrase (Flickr30K) 45K 61.52 0.5539
Noun Phrase (CC3M) 186K 62.38 0.5904
Noun Phrase (CC12M) 2.58M 62.42 0.6242
Noun Phrase (YFCC15M) 2.20M 62.45 0.6375
Noun Phrase (ALL) 5.12M 62.70 0.6498
@@ C.1 Detailed Results for All Datasets
Table 7, 8, and 9 shows all of the results on the 12 datasets omitted in Table 3, 3, and 4, respectively.